OpenStudy (anonymous):

what is the factorization of the trinomial: 3x^3+6x^2-24x a. 3x(x-2)(x-4) b. 3x(x-2)(x+4) c. 3(x-2)(x+4) d. 3(x^2-2)(x+4)

OpenStudy (anonymous):

B

OpenStudy (anonymous):

3x^3+6x^2-24x=x(3x^2+6x+-24) 3x^2+6x-24=3(x-2)(x+4) so 3x^3+6x^2-24x=3x(x-2)(x+4) answer C